# A Survey of Safe Reinforcement Learning and Constrained MDPs: A Technical Survey on Single-Agent and Multi-Agent Safety

### v1

@@ 5 Open Research Challenges and Future Directions
While significant progress has been made in SafeRL and SafeMARL, many challenges remain open. In this section, we present five research problems that, if solved, would substantially advance the field. Three of these pertain specifically to SafeMARL, reflecting the newer nature of multi-agent safety. For each problem, we describe the motivation, outline possible approaches (steps toward a solution), and reference relevant prior work to build upon.
@@
The concept of never violating constraints relates to safe exploration. Moldovan and Abbeel (2012) and others studied conditions for “safe policy learning” where certain states are absorbing traps (unsafe) and should be avoided forever. Approach like learning with a safety critic Thananjeyan and et al. (2021) or mentor-assisted exploration Saunders et al. (2018); Zhou and Li (2018); Curi et al. (2020) have been tried. However, a general solution remains elusive, especially for high-dimensional continuous tasks. Solving this problem would likely require combining learning with elements of control theory or formal methods to get the needed guarantees.
@@
In a POMDP, the agent typically maintains a belief (distribution over states). Constraints might need to be satisfied with respect to the true state (which is unknown). For instance, we might require that for all possible true states consistent with the agent’s observations, the safety constraint holds. This is a very strict condition and can be overly conservative. Alternatively, one might demand a high probability of safety given the belief.
There is work on POMDPs with chance constraints, where constraints must hold with a certain probability. Techniques often convert these into augmented state MDPs by including some memory or using scenario optimization. Another related concept is belief shielding: e.g., using human feedback to avoid ambiguous unsafe states. Solving safe RL in POMDPs could connect to robust control in partially observed systems (like robust Model Predictive Control with chance constraints). This problem remains largely open; progress would benefit fields like autonomous systems operating with imperfect sensors.
In many multi-agent applications, each agent has only partial, local observations (e.g., each car in traffic sees only nearby cars). A central authority that monitors and enforces safety for all agents may not exist. The challenge is to achieve safe multi-agent learning in a fully decentralized way: each agent makes decisions based on its local view and (optionally) limited communication, and together their behaviors ensure global safety constraints are respected.
@@
Decentralized MARL has been studied (e.g., independent learners, mean-field MARL), but safety adds extra difficulty. Zhang et al. (2024) is one of few works aiming at decentralized SafeMARL. Also relevant are distributed constrained optimization in control theory where multiple controllers ensure a global constraint (like distributed frequency control in power grids ensuring safety constraints on voltage). Techniques from graphical games or networked control systems could be applied. Success in this problem would directly impact fields like distributed robotics and network safety (e.g., ensuring no network congestion collapse via decentralized RL controllers).
@@
In competitive multi-agent scenarios, one cannot simply optimize a joint objective. Methods like CPO do not directly apply, because improving one agent’s reward might hurt another’s. We need an equilibrium concept like constrained Nash equilibrium or constrained correlated equilibrium. Another issue is that safety for one agent might depend on the behavior of others. If others act recklessly, an agent might be unable to guarantee its safety without overly sacrificing reward (or might need to assume worst-case opponents).
Constrained game equilibria have been studied in economics (e.g., Nash equilibria with budget constraints). In RL, one related field is Mean-Field Games with constraints, but results are sparse. Another is Multi-agent reinforcement learning for traffic, where multiple self-interested cars must avoid collisions (a safety constraint) – some works use hand-crafted rules or potential fields, but learning such behavior while each optimizes their own objective is largely open. If this problem is solved, it could define how autonomous systems from different stakeholders safely coexist (think of air traffic control but without a central controller—planes negotiating to avoid collisions while meeting their own goals).
@@
Non-stationarity breaks the assumptions of convergence for most RL algorithms. SafeRL adds another layer: after training, if something changes and the policy is no longer safe, the agent must detect and correct this quickly (ideally without catastrophic failure during the transition). Multi-agent adds complexity because one agent’s non-stationarity (learning or adapting) is another agent’s non-stationary environment.
Non-stationary RL is a growing area (sometimes framed as lifelong learning or non-stationary bandits). SafeRL in non-stationary settings has seen little study. One relevant angle is robust safe RL: algorithms that ensure safety under model perturbations (e.g., Chen et al. (2021); Yang et al. (2021) considered adversarial changes in cost function within limits). Another is meta-learning safety: a recent work by Grbic and Risi (2020) attempted to meta-learn a safety critic for quickly evaluating new scenarios. Achieving continual safe learning could pave the way for real-world deployment where conditions are never static.

### v2

@@ 6 Open Research Challenges and Future Directions
While significant progress has been made in SafeRL and SafeMARL, many challenges remain open. In this section, we present five research problems that, if solved, would substantially advance the field (see Figure 14). Three of these pertain specifically to SafeMARL, reflecting the newer nature of multi-agent safety. For each problem, we describe the motivation, outline possible approaches (steps toward a solution), and reference relevant prior work to build upon.
@@
The concept of never violating constraints relates to safe exploration. Moldovan and Abbeel (2012) and others studied conditions for “safe policy learning” where certain states are absorbing traps (unsafe) and should be avoided forever. Approach like learning with a safety critic mentor-assisted exploration Saunders et al. (2018); Zhou and Li (2018); Huang et al. (2018) have been tried. However, a general solution remains elusive, especially for high-dimensional continuous tasks. Solving this problem would likely require combining learning with elements of control theory or formal methods to get the needed guarantees.
@@
In a POMDP, the agent typically maintains a belief (distribution over states). Constraints might need to be satisfied with respect to the true state (which is unknown). For instance, we might require that for all possible true states consistent with the agent’s observations, the safety constraint holds. This is a very strict condition and can be overly conservative. Alternatively, one might demand a high probability of safety given the belief. There is work on POMDPs with chance constraints, where constraints must hold with a certain probability. Techniques often convert these into augmented state MDPs by including some memory or using scenario optimization. Another related concept is belief shielding: e.g., using human feedback to avoid ambiguous unsafe states. Solving safe RL in POMDPs could connect to robust control in partially observed systems (like robust Model Predictive Control with chance constraints). This problem remains largely open; progress would benefit fields like autonomous systems operating with imperfect sensors. In many multi-agent applications, each agent has only partial, local observations (e.g., each car in traffic sees only nearby cars). A central authority that monitors and enforces safety for all agents may not exist. The challenge is to achieve safe multi-agent learning in a fully decentralized way: each agent makes decisions based on its local view and (optionally) limited communication, and together their behaviors ensure global safety constraints are respected.
@@
Decentralized MARL has been studied (e.g., independent learners, mean-field MARL), but safety adds extra difficulty. Zhang et al. (2024) is one of few works aiming at decentralized SafeMARL. Also relevant are distributed constrained optimization Fioretto et al. (2018); Chang et al. (2014); Notarnicola and Notarstefano (2018); Luan et al. (2024) in control theory where multiple controllers ensure a global constraint (like distributed frequency control in power grids Parandehgheibi et al. (2016); Wang et al. (2019) ensuring safety constraints on voltage). Techniques from graphical games or networked control systems could be applied. Success in this problem would directly impact fields like distributed robotics Testa et al. (2025); Wang et al. (2022) and network safety (e.g., ensuring no network congestion collapse via decentralized RL controllers Sunassee et al. (2021); Kamau Kiarie et al. (2025)).
@@
In competitive multi-agent scenarios, one cannot simply optimize a joint objective. Methods like CPO do not directly apply, because improving one agent’s reward might hurt another’s. We need an equilibrium concept like constrained Nash equilibrium Xu et al. (2025) or constrained correlated equilibrium Boufous et al. (2024); Chen et al. (2022). Another issue is that safety for one agent might depend on the behavior of others. If others act recklessly, an agent might be unable to guarantee its safety without overly sacrificing reward (or might need to assume worst-case opponents). Constrained game equilibria have been studied in economics (e.g., Nash equilibria with budget constraints). In RL, one related field is Mean-Field Games with constraints Cannarsa et al. (2018); Barreiro-Gomez and Tembine (2019); Arjmand and Mazanti (2021); Capuani and Marigonda (2022), but results are sparse. Another is Multi-agent reinforcement learning for traffic K.J. et al. (2014); Zeynivand et al. (2022); Li et al. (2024), where multiple self-interested cars must avoid collisions (a safety constraint) – some works use hand-crafted rules or potential fields, but learning such behavior while each optimizes their own objective is largely open. If this problem is solved, it could define how autonomous systems from different stakeholders safely coexist (think of air traffic control but without a central controller—planes negotiating to avoid collisions while meeting their own goals).
@@
Non-stationarity breaks the assumptions of convergence for most RL algorithms. SafeRL adds another layer: after training, if something changes and the policy is no longer safe, the agent must detect and correct this quickly (ideally without catastrophic failure during the transition). Multi-agent adds complexity because one agent’s non-stationarity (learning or adapting) is another agent’s non-stationary environment. Non-stationary RL is a growing area (sometimes framed as lifelong learning or non-stationary bandits). SafeRL in non-stationary settings has seen little study. One relevant angle is robust safe RL: algorithms that ensure safety under model perturbations (e.g., Chen et al. (2021); Coursey et al. (2025) considered adversarial changes in cost function within limits or continual safety in non-stationary dynamics). Another is meta-learning safety: a recent work by Grbic and Risi (2020) attempted to meta-learn a safety critic for quickly evaluating new scenarios. Achieving continual safe learning could pave the way for real-world deployment where conditions are never static.
@@ 7 Conclusion
LLM Usage
ChatGPT-5.1 free version was used for polishing of initial draft text to improve the English and grammar. Claude Code was used to assist with TikZ figure creation, formatting, and bibliography management.